-6 4. Compute x5 + 3x4 + 2x3 + 1 dx.
Answer #1

Solution: 6 25+ 3x4 + 2x + 1 da d x4 = 205 24 + 3x4 2x3 24 ) dx 24 x4 6 - 2 S (x+ 3+ + 14 ) da 2 = {xdx + da x4 3 dx + 26 dx[2²+] + [3x + C2 ] + 2 [in (0) + c25+ [+4) where G, C2, C3, C4 are arbitrary constant [Sinces Sanda= s dx = ln(x) + d} where